Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 11     Entry time: 03 Feb 2024, 23:51
Author: Jörg Hallmann 
Group: MID 
Subject: Shift summary 

goals achieved:
alignment of the beam in tunnel and hutches including refinement of the beam in the center of rotation 
check of cameras and optical components
initial check of the mirror chamber (successfully)

issues:
smarACT motors in the DES tricky - some are in error state and the filter wheel is moving and does not stop - positioned it and removed afterwards the cable - EEE intervention next week necessary
camera of Pop-in imager II45-2 in XTD6 went into error - DOC solved the issue

ELOG V3.1.4-7c3fd00